OpenStudy (anonymous):

what is the soultion of the system?-4x+3y=-12

OpenStudy (anonymous):

trying to solve for y?

OpenStudy (anonymous):

x and y @sleung

OpenStudy (anonymous):

You only have one equation, so there are infinite solutions.

OpenStudy (anonymous):

\[-4x=-12-3y\] \[x=3+\frac{ 3 }{ 4 }y\] \[3y=-12+4x\] \[3y=12+4x\] \[y=4+\frac{ 4 }{ 3 }x\]